OpenStudy (johnjules2000):

Solving equations with variables on both sides. Must show work five x minus two plus three x == seventeen plus twelve x minus 23 5x-2+3 = 17+12x-23

OpenStudy (ja1):

Simple, let's follow first combine like terms,can you do that?

OpenStudy (johnjules2000):

somewhat ish

OpenStudy (ja1):

Ok well first start off by telling me what the like terms are.

OpenStudy (johnjules2000):

5x 3x and 12x

OpenStudy (ja1):

Perfect, so first let's combine the ones on the left side, 5x and 3x. Remember to carry the sign!

OpenStudy (ja1):

So 5x + 3x = ?

OpenStudy (johnjules2000):

8x

OpenStudy (ja1):

Excellent! So know I always like to think of it this way "Take the baby to the mommy", in other words take the smaller sum, in this case 8x, and combine it with the bigger sum, 12x. Remember to do the inverse operation :D

OpenStudy (johnjules2000):

so is it 12x - 8x?

OpenStudy (ja1):

Ok sorry I was afk, anyways yes since 8x is positive.

OpenStudy (johnjules2000):

so now the equation is 2= 4x+ 17 - 23?

OpenStudy (ja1):

So then afterwards you should get: \[-2 = 4x + 17 - 23 \]

OpenStudy (johnjules2000):

ok

OpenStudy (ja1):

Precisely, so now you would combine the other like terms, 17 and -23. P.S. On that -2 remember to carry the sign with it :)

OpenStudy (johnjules2000):

and 17- 23 = -6 so then the equation should be -2= 4x + -6

OpenStudy (johnjules2000):

i think i got it ......... x= 1???

OpenStudy (ja1):

Yes! you would basically do: -2 = 4x - 6 +6 +6 4 = 4x /4 /4 x = 1

OpenStudy (ja1):

Well done :D

OpenStudy (ja1):

My only tip is to always remember those signs, they can get tricky :)

OpenStudy (johnjules2000):

thank you will you also help me with this one 4 ( x + 5) + 2 = x + 3

OpenStudy (ja1):

Ok, in this case we need to get those pesky parenthesis out of the way, can you show me the distributed form?

OpenStudy (johnjules2000):

4x + 20 + 2 = x + 3

OpenStudy (ja1):

Yes! So now do exactly the same as last time, combine the 4x and x (1x)

OpenStudy (johnjules2000):

-3 x ?

OpenStudy (ja1):

Well technically you are right, but remember, it's easier to take the smaller sum to the larger one, so for this example let's do 4x - 1x = 3x

OpenStudy (johnjules2000):

so now I have 20 + 2 = -3x + 3 right?

OpenStudy (ja1):

Eh scratch what I said, let's just move on :) So now let's add the 20 and the 2

OpenStudy (johnjules2000):

22 = 3x + 3 ? right ?

OpenStudy (ja1):

Yes, so now let's combine that 3 and the 22 P.S. Again, remember the negative sign on that 3 :)

OpenStudy (johnjules2000):

19

OpenStudy (ja1):

Wait I think we made a mistake somewhere, one second...

OpenStudy (ja1):

are you sure you copied the problem correctly?

OpenStudy (johnjules2000):

yes

OpenStudy (ja1):

Hmmm, I am getting a decimal here, is that something your teacher has had you do before?

OpenStudy (johnjules2000):

no but she said we will get a decimal answer

OpenStudy (ja1):

Oh ok, I found it weird since most teachers don't like to give decimal problems, anyway so let's continue, 3x = 19 so we divide 19 by 3

OpenStudy (johnjules2000):

i got it

OpenStudy (ja1):

Ok glad I could help! A medal would be much appreciated :)